OpenStudy (anonymous):

Will fan and medal!!!!! The length of each of the 2 congruent sides of an isosceles triangle is 4x+2. The length of the third side is 2x The perimeter is 29 centimeters. What is an equation that could be used to find the value of x. Solve for x and find the length of all three sides.

jhonyy9 (jhonyy9):

because do you know that the perimeter of an isoscele triangle is formed by adding the 2 equal lenght side with to the 3rd side length - yes ? so you know that the length of these two congruent sides is 4x+2 and the 3rd side has a length of 2x how will get the perimeter ?

OpenStudy (anonymous):

i dont know!

jhonyy9 (jhonyy9):

P=two equale sides + 3rd side P=2*(4x+2) +2x 29=2*(4x+2)+2x yes ? so can you calcule the value of x from this equation ?

OpenStudy (anonymous):

yeah, it'll be 5/2 or 2 1/2

jhonyy9 (jhonyy9):

29=8x+4+2x 25=10x x=25/10 =2,5 cm

jhonyy9 (jhonyy9):

so now you need to calcule the realy length of sides - how many will them be - ok. ?

jhonyy9 (jhonyy9):

just substitute the value of x inside this length of sides 4x+2 and 2x

OpenStudy (anonymous):

ok

jhonyy9 (jhonyy9):

these two equale sides will have 4*2,5 +2 = ? cm and the 3rd will be 2*2,5 cm right ?

jhonyy9 (jhonyy9):

than you can ending it bye bye and was my pleasure anytime than i m here online

OpenStudy (anonymous):

Thank you so much!!!

jhonyy9 (jhonyy9):

yw good luck bye bye